Two two cars are at a distance of 100 km from each other .Consider the first car to be ,A and the second as B.Now these cars are travelling at different speeds.If these cars travel in the same direction , these meet in 5 hour and if these travel in directions opposite to each other's, they meet in 1 hour .Find the speed of the two cars in km/ h.

Grade:8

1 Answers

Nishant Vora IIT Patna
7 years ago
relative speed = relative distance / time taken

Let A’s speed be x and B’s speed be y

so, for same direction, rel. speed is x-y
for opp dirn, rel. speed is x+y

hence, x – y = 100 / 5 ----------(1)
and x + y = 100 / 1 …..........(2)

Solve eqn 1 and 2 now,

hence x = 60 and y = 40